# Enhancing Redundancy-based Automated Program Repair by Fine-grained Pattern Mining

### v1

@@ Abstract.
Redundancy-based automated program repair (APR), which generates patches by referencing existing source code, has gained much attention since they are effective in repairing real-world bugs with good interpretability. However, since existing approaches either demand the existence of multi-line similar code or randomly reference existing code, they can only repair a small number of bugs with many incorrect patches, hindering their wide application in practice. In this work, we aim to improve the effectiveness of redundancy-based APR by exploring more effective source code reuse methods for improving the number of correct patches and reducing incorrect patches. Specifically, we have proposed a new repair technique named Repatt, which incorporates a two-level pattern mining process for guiding effective patch generation (i.e., token and expression levels). We have conducted an extensive experiment on the widely-used Defects4J benchmark and compared Repatt with eight state-of-the-art APR approaches. The results show that our approach complements existing approaches by repairing 15 unique bugs compared with the latest deep learning-based methods and 19 unique bugs compared with traditional repair methods when providing the perfect fault localization. In addition, when the perfect fault localization is unknown in real practice, Repatt significantly outperforms the baseline approaches by achieving much higher patch precision, i.e., 83.8%. Moreover, we further proposed an effective patch ranking strategy for combining the strength of Repatt and the baseline methods. The result shows that it repairs 124 bugs when only considering the Top-1 patches and improves the best-performing repair method by repairing 39 more bugs. The results demonstrate the effectiveness of our approach for practical use.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; December 2023; price: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Automated program repair (APR) techniques have the potential to significantly reduce the debugging overhead of human developers and thus have been widely studied in the last decades. To date, a large number of APR approaches have been proposed and developed (Wen et al., 2018; Xiong et al., 2017; Xuan et al., 2017; Liu et al., 2019; Tan et al., 2016; Weimer et al., 2009; Long and Rinard, 2016; Jiang et al., 2018; Xin and Reiss, 2017b; Nguyen et al., 2013; B. Le et al., 2016). A typical APR approach takes a faulty program and a set of test cases, where at least one triggers the bug in the program, as inputs, and produces one or more (if possible) plausible patches that can make all the test cases pass. The repair process can be typically viewed as a search problem, where the search space is all possible programs. Therefore, the core challenge for APR techniques is how to effectively and efficiently locate the desired patches within limited computing resources and time budget. To facilitate this process, existing techniques have employed multiple data sources, e.g., similar code (Jiang et al., 2018; Xin and Reiss, 2017b), historical patches (Le et al., 2016, 2017), repair templates (Liu et al., 2019; Kim et al., 2013; Saha et al., 2017), etc., and leveraged diverse search algorithms, e.g., random search (Qi et al., 2014), genetic programming (Weimer et al., 2009; Le Goues et al., 2012), and advanced deep learning (Zhu et al., 2021; Ye et al., 2022b; Jiang et al., 2023a, 2021), aiming to effectively refine the search space and thus improve the repair ability and efficiency.
@@
Among existing APR techniques, redundancy-based (Chen, 2018; White et al., 2019; Yang, 2021; Chen and Monperrus, 2018) techniques have attracted much attention and have gained big success. These techniques are on the basis of the plastic surgery hypothesis that “changes to a codebase contain snippets that already exist in the codebase at the time of the change, and these snippets can be efficiently found and exploited (Barr et al., 2014)”. Therefore, redundancy-based APR approaches generate patches by searching and referencing existing code. For example, GenProg (Weimer et al., 2009; Le Goues et al., 2012), the pioneer of modern APR techniques, generates patches by directly reusing existing code under the guidance of genetic programming algorithms. Although the recent advance of large-scale pre-trained models has shown the promise to repair more bugs (Zhu et al., 2021; Jiang et al., 2023a; Ye et al., 2022b), they still suffer from the interpretability issue due to the complexity and the random nature of deep neural networks, and mainly repair simple bugs (Ye et al., 2022b). As reported by previous studies (Liu et al., 2020; Motwani et al., 2022; Wang et al., 2019), the patches generated by SimFix, a representative redundancy-based APR approach, are of high quality under various testing scenarios. In addition, redundancy-based techniques also complement those learning-based techniques (Zhu et al., 2021; Ye et al., 2022b; Jiang et al., 2023a) (Our evaluation results in Section 5 further prove it). Additionally, redundancy-based techniques tend to have better interpretability as they are on the basis of the plastic surgery hypothesis, and the patches are typically generated by referring human-written code snippets. In particular, they can repair a portion of relatively complex bugs (Yang et al., 2023; Jiang et al., 2018; Xin and Reiss, 2017b).
@@
Although existing redundancy-based techniques are demonstrated to be effective for repairing real-world bugs, they still face two major challenges: (1) How to effectively and efficiently locate the reference code elements among a large-scale codebase? Existing approaches identify similar code snippets as the references for patch generation by leveraging code features like structures (Jiang et al., 2018; Yang et al., 2023) and token sequences (Xin and Reiss, 2017b). The idea behind them is that similar code snippets have a higher possibility to implement the same functionality and thus can guide bug fixes. However, as it will be shown in our preliminary study (Section 2.1), the coarse-grained similar code snippets for patch generation are indeed rare in real practice. The limited number of similar code snippets hinders them from repairing a large number of bugs and thus decreases their usability. (2) How to reuse existing code for constructing new patches? Existing approaches reuse code by comparing the difference between faulty and reference code via abstract syntax trees (Jiang et al., 2018; Xin and Reiss, 2017b) or program dependency graphs (Yang et al., 2023). However, code elements under different contexts tend to vary greatly in structures (Rattan et al., 2013), limiting the effectiveness of existing techniques, i.e., repairing a small number of bugs with many incorrect patches.
@@
As explained above, existing approaches reuse code either completely randomly (e.g., RSRepair (Qi et al., 2014) and GenProg (Weimer et al., 2009)) or depending on the existence of similar code snippets of multiple lines (e.g., SimFix (Jiang et al., 2018) and TransplantFix (Yang et al., 2023)). As it will be presented in our preliminary study, though the reusable code elements for patch generation may appear at different granularities, most of them actually exist at the finer-grained level, e.g., 89.3% reusable code elements contain less than three tokens (Section 2.1). In this paper, we investigate the possibility of improving redundancy-based APR techniques by reusing code elements at a finer-grained level. Specifically, the basic idea of our approach is that fine-grained code elements related to similar code semantics tend to co-appear nearby in the program, which has been well-studied by existing research (Hindle et al., 2012; Ray et al., 2016; Khanfir et al., 2022; Tu et al., 2014). We call this phenomenon locality property of source code. Based on this property, we proposed a new APR approach that effectively guides patch generation by identifying the usage patterns of fine-grained code elements for confining the patch space and thus overcomes the first challenge of redundancy-based APR. To overcome the second challenge, i.e., reusing existing code under different contexts, we designed a new code representation method, i.e., S-TAC, which decomposes complex code expressions and statements into a unified form by ignoring context-specific features (e.g., code structures and operators). It can reduce the negative impact of code structures on code matching during patch generation but still preserve the locality property of source code.
@@
To evaluate the performance of our approach, we have implemented it as an APR tool named Repatt and conducted an extensive study by comparing it with eight state-of-the-art APR approaches, including three best-performing traditional repair techniques and five latest deep learning-based techniques. The experimental results show that our approach complements existing approaches by correctly repairing many unique bugs. Specifically, Repatt repairs 19 and 15 unique bugs that cannot be repaired by the traditional and deep learning-based methods, respectively. In particular, 9 bugs repaired by Repatt have never been repaired by all the baselines. Additionally, when the perfect fault localization is unknown, which is a more realistic scenario, Repatt can significantly outperform the baseline approaches by achieving 33.8%-51.7% higher patch precision. Additionally, we also made the first attempt to combine the strength of Repatt and multiple existing approaches by further designing a patch ranking strategy. The evaluation result shows that the combined method can effectively repair 124 bugs when only considering Top-1 patches, 39 more bugs than the best-performing method. Our results demonstrate that our approach is indeed effective in improving the performance of existing APR techniques.
@@
  • A new redundancy-based APR technique that can flexibly reuse source code at finer-grained granularities.
@@
  • We make all our experimental results and implementations publicly available to facilitate replication and comparison.
The remaining of the paper is organized as follows. Section 2 introduces our preliminary study and running examples to motivate our approach. Section 3 introduces the details of our approach. Then, Section 4 presents the experimental setup while Section 5 analyzes the evaluation results. We discuss the threats to validity and related work in Section 6 and 7, respectively. Finally, Section 8 concludes the paper.
2. Motivation
2.1. Preliminary Study
@@
Figure 1. The length distribution of code elements that can be found in the faulty program for repairing related bugs.
Although many redundancy-based APR techniques have been proposed and produced promising results (Jiang et al., 2018; Xin and Reiss, 2017b; Liu et al., 2020; Motwani et al., 2022; Wang et al., 2019), these techniques reuse existing code elements by either random search without considering the relevance of reference code (Weimer et al., 2009; Le Goues et al., 2012; Weimer et al., 2013; Qi et al., 2014) or searching similar code at a coarse-grained granularity, e.g., code blocks (Jiang et al., 2018; Xin and Reiss, 2017b). However, no study explores the potential of searching and reusing finer-grained code elements, e.g., tokens, for patch generation. To investigate the feasibility and necessity of reusing finer-grained code elements for APR, we have conducted a preliminary study by following the idea of Barr et al. (Barr et al., 2014). Specifically, the objective of our study is to understand the granularity of reusable code elements for patch generation. To achieve that, we performed a greedy search process. When given a bug and its associated patch, we first extracted the newly added code elements in the patch and searched for them in the faulty program from coarse-grained to finer-grained. That is, if a code element (e.g., “a + f(b,c)”) was not found, we would further decompose it into finer-grained code elements based on the syntax tree structure of the code (e.g., “a”, “+”, and “f(b,c)”). This process was iteratively proceeded until the elements cannot be further decomposed into finer-grained ones (e.g., variables). Finally, we analyzed the granularity of code elements that can be found in faulty programs. In particular, we use the number of tokens to measure the granularity of code elements since code structures in the Abstract Syntax Tree (AST) cannot explicitly reveal the granularity of code elements. For example, an expression in the AST can represent either a simple variable like “a” or a complex expression like “f(a, b)”. We conducted our empirical study on two widely-used datasets of real-world bugs, i.e., Defects4J (Just et al., 2014) and Bugs.jar (Saha et al., 2018). In particular, we filtered bugs whose patches involve more than one Java file since they are still hard to repair at present (White et al., 2019; Jiang et al., 2018; Xiong et al., 2017; Xia et al., 2023; Jiang et al., 2023a; Zhu et al., 2021; Chen et al., 2019; Li et al., 2020). As a result, 564 bugs in Defects4J and 609 bugs in Bugs.jar from 24 diverse projects were included. Figure 1 presents the statistical results. The x-axis denotes the number of tokens in a code element and the y-axis denotes the percentages of code elements that can be reused for patch generation. Please note that a patch may involve multiple code elements. The results show on average more than 89.3% code elements contain less than three tokens, and 42% elements only contain one token. In contrast, only a very limited portion of code elements exist at a coarse-grained granularity (i.e., containing more than three tokens). In particular, this is not a special case but holds over different programs in our study. The result indicates that the reusable code elements indeed exist at a fine-grained granularity, which will significantly limit the effectiveness of existing redundancy-based APR approaches. Moreover, it also reflects that designing APR approaches by reusing finer-grained code elements is promising. However, achieving accurate and fine-grained code reuse for patch generation is challenging. The reasons are twofold: (1) It will significantly enlarge the search space of candidate patches, and thus make it harder to correctly and efficiently find the desired code elements. (2) Plausible (i.e., can pass all the test cases) but incorrect patches will be more easily generated due to the problem of weak test suite (Qi et al., 2015; Xin and Reiss, 2017a; Xiong et al., 2018). It will cause heavier manual effort for patch validation and thus significantly affect the usability of the APR approach in practice. To overcome this challenge, in this work, we propose an effective APR approach that can efficiently locate the desired code elements for patch generation. The basic idea of our approach is that code elements related to similar code semantics tend to co-appear nearby in the programs. For example, the variables “numerator” and “denominator” tend to be used together. We call this phenomenon the locality property of source code. Consequently, it would be possible to employ those code elements, which tend to co-appear with the ones that already exist in the faulty location, for patch generation. As it will be presented in our evaluation, the locality property does help to repair real-world bugs. The details will be introduced in Section 3.
2.2. Running Example
@@
Motivated by our preliminary study, reusing finer-grained code elements for patch generation and boosting the effectiveness of redundancy-based program repair is feasible and promising. In this section, we employ running examples of real-world bugs to further explain how the locality property of source code can be leveraged to guide patch generation. Listing 1 presents the patch code of Codec-3 from the Defects4J (Just et al., 2014), where a line starting with “+” denotes newly added code while starting with “-” denotes code to be deleted.
452   if ((contains(value, 0 ,4, "VAN ", "VON ") || ...)){
@@
In this example, the constant parameter “4” was mistakenly used at line 455, and the desired parameter is “3”. This bug cannot be fixed by existing redundancy-based APR techniques due to either the large search space (e.g., GenProg (Weimer et al., 2009; Le Goues et al., 2012), RSRepair (Qi et al., 2014), etc.) or the inexistence of multi-line similar code snippets for reference (e.g., SimFix (Jiang et al., 2018), transplantFix (Yang et al., 2023), etc.). However, when it comes to the fine-grained token level, we will find that the constant number “3” usually co-appears with the string “IER” elsewhere in the program since the former really denotes the length of the latter string. In other words, these two tokens are strongly correlated to each other (i.e., the locality property of source code). As a result, it will be feasible and easier to effectively find the desired code elements for patch generation under the guidance of this property. In fact, the locality property of source code may not only exist at the token level but may also exist at higher levels, e.g., expression or statement levels. For example, Listing 2 presents the patch code of Gson-17 from the Defects4J benchmark, while Listing 3 presents a reference code that can be used for patch generation. According to the reference code, the expression “in.peek()” may be correlated with the expressions “JsonToken.NULL”, “in.nextNull()” and “return null”. In other words, the reference code can provide the required code elements for generating the desired patch in this example. However, it also can be seen that although these expressions can be taken for patch generation, how to utilize them is still challenging since the possible combinations of code changes according to the reference code are still too many, such as replacing the throw statement with “in.nextNull()” or inserting “in.nextNull”. In addition, replacing the conditional expression also does not fix the bug due to the incorrect operator “!=”. In particular, the code structures of reference code and the faulty code may also be different in practice, making patch generation harder.
@@
To address this challenge, we propose a novel code representation method that decomposes complex expressions into a unified simple form, which ignores operators and complex code structures, e.g., !=, for, while, etc., but still preserves the locality property of source code. For example, the transformed representation of the faulty code and the reference code shown in Listings 2 and 3 are presented in Figure 4. As a result, we can generate candidate patches by aligning these two representations.
3. Framework
Figure 2. Overview of our approach Repatt
In this section, we introduce the details of our approach (named Repatt). Figure 2 presents the overview of Repatt. In general, Repatt generates candidate patches by mining the usage patterns of code elements at different granularities as guidance based on the plastic hypothesis and the locality property of source code. Particularly, in order to repair bugs that demand code changes at different granularities, Repatt comprises two components: token-level pattern mining and expression-level code search. The former aims to repair bugs that demand code changes within an individual line (ref. Listing 1), while the latter aims to repair bugs that demand code changes across multiple code lines (ref. Listing 2).
Specifically, the token-level pattern mining is an offline process whose target is to construct a code pattern database that can be effectively and efficiently queried in the subsequent online repair process. On the contrary, the expression-level code search is an online process that aims to search for a set of code elements that can be referenced for patch generation. The reasons for their different search strategies are twofold: (1) As aforementioned, token-level pattern mining aims to repair single-line bugs. Therefore, we can only consider the usage patterns of tokens co-appearing within single lines, which can be small-scale. In contrast, the expression-level code usage patterns will tend to produce a much larger search space when considering all possible combinations of expressions from different lines of code. (2) Token-level patterns tend to produce many small code changes and thus may severely affect the efficiency of the repair process, while the expression-level patterns correlated to a certain location (i.e., the faulty code) will be very limited according to the statistical results in our study (refer to Figure 1). Therefore, we confine the token-level patterns to those that are frequent. In other words, we need to analyze whether the two patterns are the same or not for tokens.
3.1. Offline Pattern Mining

### v3

# Boosting Redundancy-based Automated Program Repair by Fine-grained Pattern Mining

@@ Abstract
Redundancy-based automated program repair (APR), which generates patches by referencing existing source code, has gained much attention since they are effective in repairing real-world bugs with good interpretability. However, since existing approaches either demand the existence of multi-line similar code or randomly reference existing code, they can only repair a small number of bugs with many incorrect patches, hindering their wide application in practice. In this work, we aim to improve the effectiveness of redundancy-based APRs by exploring more effective source code reuse methods for improving the number of correct patches and reducing incorrect patches. Specifically, we have proposed a new repair technique named Repatt, which incorporates a two-level pattern mining process for guiding effective patch generation (i.e., token and expression levels). We have conducted an extensive experiment on the widely-used Defects4J benchmark and compared Repatt with ten state-of-the-art APR approaches. The results show that it complements existing approaches by repairing 9 unique bugs compared with the latest Large Language Model (LLM)-based and deep learning-based methods and 19 unique bugs compared with traditional repair methods when providing the perfect fault localization. In addition, when the perfect fault localization is unknown in real practice, Repatt significantly outperforms the baseline approaches by achieving much higher patch precision, i.e., 83.8%, although it repairs fewer bugs. Moreover, we further proposed an effective patch ranking strategy for combining the strength of Repatt and the baseline methods. The result shows that it repairs 124 bugs when only considering the Top-1 patches and improves the best-performing repair method by repairing 39 more bugs. The results demonstrate the effectiveness of our approach for practical use.
Index Terms:
Automated program repair, Pattern mining, Program debugging
I Introduction
@@
Automated program repair (APR) techniques have the potential to significantly reduce the debugging overhead of human developers and thus have been widely studied in the last decades. To date, a large number of APR approaches have been proposed and developed [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. A typical APR approach takes a faulty program and a set of test cases, where at least one triggers the bug in the program, as inputs, and produces one or more (if possible) plausible patches that can make all the test cases pass. The repair process can be typically viewed as a search problem, where the search space is all possible programs. Therefore, the core challenge for APR techniques is how to effectively and efficiently locate the desired patches within limited computing resources and time budget. To facilitate this process, existing techniques have employed multiple data sources, e.g., similar code [8, 9], historical patches [12, 13], repair templates [4, 14, 15], etc., and leveraged diverse search algorithms, e.g., random search [16], genetic programming [6, 17], and advanced Large Language Model (LLM) and deep learning [18, 19, 20, 21, 22, 23, 24], aiming to effectively refine the search space and thus improve the repair ability and efficiency.
@@
Among existing APR techniques, redundancy-based [25, 26, 27, 28] techniques have attracted much attention and achieved considerable success. These techniques are on the basis of the plastic surgery hypothesis that “changes to a codebase contain snippets that already exist in the codebase at the time of the change, and these snippets can be efficiently found and exploited [29]”. In other words, the code snippets for patch generation can be found within the buggy projects themselves. Therefore, redundancy-based APR approaches leverage the plastic surgery hypothesis to generate patches by searching and referencing existing code. For example, GenProg [6, 17], the pioneer of modern APR techniques, generates patches by directly reusing existing code under the guidance of genetic programming algorithms. Although the recent advance of LLMs has shown the promise to repair more bugs [20, 22, 23, 24], they still suffer from the interpretability issue due to the complexity and the random nature of deep neural networks as they are typically used in a black-box fashion. As reported by previous studies [30, 31, 32], the patches generated by SimFix, a representative redundancy-based APR approach, are of high quality under various testing scenarios. In addition, redundancy-based techniques also complement those learning-based techniques [18, 19, 20] (our evaluation results in Section V further prove it). Additionally, redundancy-based techniques tend to have better interpretability as they are on the basis of the plastic surgery hypothesis, and the patches are typically generated by referring human-written code snippets. In particular, they can repair a portion of relatively complex bugs [33, 8, 9].
@@
Although existing redundancy-based techniques are demonstrated to be effective for repairing real-world bugs, they still face two major challenges: (1) How to effectively and efficiently locate the reference code elements among a large-scale codebase? Existing approaches identify similar snippets using structural [8, 33] or token-based [9] features, assuming similarity implies functional equivalence. However, as shown in our preliminary study (Section II-A), coarse-grained similar snippets are rare in practice, limiting their usability. (2) How to reuse existing code for constructing new patches? Existing approaches reuse code by comparing the difference between faulty and reference code via abstract syntax trees [8, 9] or program dependency graphs [33]. However, code elements under different contexts tend to vary greatly in structures [34], limiting the effectiveness of existing techniques, i.e., repairing a small number of real-world bugs with many incorrect patches.
@@
As explained above, existing redundancy-based approaches reuse code either completely randomly (e.g., RSRepair [16] and GenProg [6]) or depending on the existence of similar code snippets of multiple lines (e.g., HDRepair [35], SimFix [8] and TransplantFix [33]). However, our preliminary study reveals that most reusable elements appear at a finer-grained level—e.g., 89.3% contain fewer than three tokens (Section II-A). This suggests that finer-grained code reuse is crucial. In this paper, we investigate the possibility of improving redundancy-based APR techniques by reusing code elements at a finer-grained level. Specifically, the basic idea of our approach is that fine-grained code elements related to similar code semantics tend to co-appear nearby in the program, which has been well-studied by existing research [36, 37, 38, 39]. This phenomenon, which we refer to as the locality property of source code, suggests that semantically related code components—such as variables, functions, or expressions—are often found in close proximity within the same code block or module. Based on this property, we proposed a new APR approach that effectively guides patch generation by identifying the usage patterns of fine-grained code elements for confining the patch space and thus overcomes the first challenge of redundancy-based APR. To overcome the second challenge, i.e., reusing existing code under different contexts, we designed a new code representation method, i.e., S-TAC, which decomposes complex code expressions and statements into a unified form by ignoring context-specific features (e.g., code structures and operators). It can reduce the negative impact of code structures on code matching during patch generation but still preserve the locality property of source code.
@@
To evaluate the performance of our approach, we have implemented it as an APR tool named Repatt and conducted an extensive study by comparing it with ten state-of-the-art APR approaches, including four best-performing traditional repair techniques and six latest LLM-based and deep learning-based techniques. The experimental results show that although our approach did not outperform all baselines considering all comparison aspects, it complements existing approaches by correctly repairing many unique bugs. Specifically, Repatt repairs 19 unique bugs that cannot be repaired by traditional methods and 9 unique bugs that neither LLM-based nor deep learning-based methods can repair. In particular, 5 bugs repaired by Repatt have never been repaired by all the baselines. Additionally, when the perfect fault localization is unknown, which is a more realistic scenario, Repatt can significantly outperform the baseline approaches by achieving 15.6%-51.7% higher patch precision, which is critical for practical use of APRs. Additionally, we also made the first attempt to combine the strength of Repatt and multiple existing approaches by further designing a patch ranking strategy. The evaluation result shows that the combined method can effectively repair 124 bugs when only considering Top-1 patches, 39 more bugs than the best-performing method (i.e., TBar). Our results demonstrate that our approach is indeed effective in improving the performance of existing APR techniques.
@@
  • A new redundancy-based APR technique that can flexibly reuse source code at a finer granularities.
@@
  • We make all our experimental results and implementations publicly available to facilitate replication and comparison [40].
II Motivation
II-A Preliminary Study
@@
Figure 1: The length distribution of code elements that can be found in the faulty program for repairing 564 and 609 bugs from benchmarks Defects4J and Bugs.jar.
As previously mentioned, no existing studies have explored the potential of searching for and reusing finer-grained code elements for patch generation. To investigate the feasibility and necessity of this approach in APR, we conducted a preliminary study following the idea proposed by Barr et al. [29].
Specifically, our study aims to understand the granularity of reusable code elements for patch generation. Given a bug and its associated patch, we first extracted newly added code elements from the patch and searched for them in the faulty program, starting from coarse-grained elements and gradually decomposing them into finer-grained components based on the syntax tree structure. That is, if a code element (e.g., “a + f(b,c)”) was not found, we would further decompose it into finer-grained code elements of the code (e.g., “a”, “+”, and “f(b,c)”). This process iterated until no further decomposition was possible. Finally, we analyzed the granularity of code elements that can be found in faulty programs. In particular, since AST structures cannot directly reflect granularity (e.g., an expression in the AST can be either a variable like “a” or a complex expression like “f(a, b)”.), we measured granularity by the number of tokens in each code element.
We conducted our empirical study on two widely-used datasets of real-world bugs, i.e., Defects4J [41] and Bugs.jar [42]. In particular, we filtered out bugs whose patches involve more than one Java file since they are still hard to repair at present [26, 8, 2, 43, 20, 18, 44, 45], resulting in 564 bugs from Defects4J and 609 bugs from Bugs.jar across 24 projects. Figure 1 presents the statistical results. The x-axis denotes the number of tokens in a code element and the y-axis denotes the percentages of code elements that can be reused for patch generation. The results show on average more than 89.3% code elements contain less than three tokens, and 42% elements only contain one token. In contrast, only a very limited portion of code elements exist at a coarse-grained granularity (i.e., more than three tokens). In particular, this is not a special case but holds over different programs in our study. The result indicates that the reusable code elements indeed exist at a fine-grained granularity, which will significantly limit the effectiveness of existing redundancy-based APR approaches. Moreover, it also reflects that designing APR approaches by reusing finer-grained code elements is promising.
However, achieving accurate and fine-grained code reuse for patch generation is challenging. The reasons are twofold: (1) It will significantly enlarge the search space of candidate patches, making it harder to efficiently and correctly find the desired code elements. (2) Due to weak test suites [46, 47, 48], plausible (i.e., can pass all the test cases) but incorrect patches are more likely to be generated, increasing the manual effort required for validation and significantly affect the practical usability of APR approaches. To overcome these challenges, we propose an effective APR approach that efficiently locates the desired code elements for patch generation leveraging the locality property of source code as its core idea. The details will be introduced in Section III.
II-B Running Example
@@
In this section, we use real-world bug examples to illustrate how the locality property of source code can guide patch generation. Listing 1 presents the patch code of Codec-3 from the Defects4J [41], where a line starting with “+” denotes newly added code while starting with “-” denotes deleted code.
In this example, the constant parameter “4” was mistakenly used at line 455, and the desired parameter is “3”. This bug cannot be fixed by existing redundancy-based APR techniques due to either the large search space (e.g., GenProg [6, 17], RSRepair [16], etc.) or the nonexistence of multi-line similar code snippets for reference (e.g., SimFix [8], TransplantFix [33], etc.). However, when it comes to the fine-grained token level, we will find that the constant number “3” usually co-appears with the tokens “value” and “index” elsewhere in the program. These tokens exhibit a strong correlation (i.e., the locality property of source code). Leveraging this property makes it both feasible and efficient to locate the correct code elements for patch generation.
452 if ((contains(value, 0 ,4, "VAN ", "VON ") $|$ ...)){
@@
In fact, the locality property of source code may not only exist at the token level but may also exist at higher levels, e.g., expression or statement levels. For example, Listing 2 presents the patch code of Gson-17 from the Defects4J benchmark, while Listing 3 presents a reference code that can be used for patch generation. According to the reference code, the expression “in.peek()” may be correlated with the expressions “JsonToken.NULL”, “in.nextNull()” and “return null”. It can be seen that, although the reference code can provide the required code elements for generating the desired patch in these examples, how to utilize them is still challenging since the possible combinations of code changes according to the reference code are still too many, such as replacing the throw statement with “in.nextNull()” or inserting “in.nextNull”. In addition, replacing the conditional expression also does not fix the bug due to the incorrect operator “!=”. In particular, the code structures of reference code and the faulty code may also be different in practice, making patch generation harder. To address this challenge, we propose a novel code representation method that decomposes complex expressions into a unified simple form, which ignores operators and complex code structures, e.g., !=, for, while, etc., but still preserves the locality property of source code.
III Framework
Figure 2: Overview of our approach Repatt. Given a buggy project, ❶ Repatt first mines a set of token sequence patterns and represents them as prefix trees. ❷ Then, based on the faulty code snippet, Repatt online searches a set of similar code snippets and transforms them into S-TAC form as reference patterns. ❸ Finally, Repatt generates candidate patches by matching the faulty code with the reference patterns and validates them one by one.
In this section, we introduce the details of our approach (named Repatt). Figure 2 presents the overview of Repatt. In general, Repatt generates candidate patches by leveraging the plastic surgery hypothesis and the locality property of source code, utilizing both token-level pattern mining and expression-level code search to handle different granularities of code changes. Token-level pattern mining, an offline process, constructs a query-efficient code pattern database for repairing single-line bugs (ref. Listing 1). In contrast, expression-level code search, an online process, identifies reference code elements for multi-line fixes (ref. Listing 2). Their differing strategies stem from (1) token-level mining focusing on small-scale patterns within single lines, while expression-level patterns span multiple lines, significantly expanding the search space, and (2) token-level patterns often introducing numerous small changes, impacting efficiency, while the expression-level patterns correlated to a certain location (i.e., the faulty code) will be very limited (refer to Figure 1). Therefore, we confine the token-level patterns to those that are frequent to balance efficiency.
III-A Offline Pattern Mining
Based on our preliminary study (Section II-A), most reusable code elements for patch generation exist at a fine-grained level, typically one or two tokens. To leverage these elements, Repatt employs a novel token-level pattern mining algorithm guided by the locality property of source code. This process is conducted offline over the given faulty project, constructing a database of token usage patterns to support efficient online patch generation. As aforementioned, we only consider the token patterns within single lines to confine the search space and avoid involving too much noise. In particular, it is common that the tokens used in different code lines share both commonalities and diversities. For example, the token sequence in some code lines can be (“a”, “b”, “c”), while it may also be (“a”, “d”, “c”) in some other code lines. To preserve the semantics of programs, we keep the token orders and design a skip-fashion pattern mining algorithm to address the problem of diversity among code lines. Additionally, each token is assigned a unique ID for efficient comparison, while separators (e.g., “,” and “(”) and structural keywords (e.g., “if” and “for”) are removed as they do not contribute to token-level patch generation.
Definition 1
(Prefix Embedding Tree): a prefix embedding tree constitutes a set of nodes, each of which is a tuple of $t=\left\langle tok,id,p,C,sup\right\rangle$, where $id$ denotes a unique embedding of a certain token toktok, $p$ denotes the parent node of $t$, $C$ denotes a set of child nodes of $t$, while supsup represents the frequency of the token sequence from the root to the current node $t$.
Furthermore, we use $t.childNode(id)$ to obtain the child node of $t$ with the embedding $id$, and use $t.setChildNode(id,chd)$ to set the node chdchd with embedding id as the child node of $t$. In addition, we use findOrCreate(trees, tok) to find the tree rooted at the token tok in trees. According to these notions, we present the pattern mining process in Algorithm 1. In general, given the faulty program, Repatt first decomposes all the code lines in the program into a set of token sequences (i.e., ESs), where each token sequence corresponds to one line of code. Then, it constructs the prefix embedding trees (i.e., trees), which stores the token patterns with corresponding usage frequencies.
1